# Seer: Online Context Learning for Fast Synchronous LLM Reinforcement Learning

### v1

@@ 1 Introduction
To improve hardware utilization, recent works have explored asynchronous rollout systems [zhong2025streamrl, fu2025areal, han2025asyncflow, sheng2025laminar], which overlap the rollout and training phases. While this approach can reduce end-to-end iteration time, it comes at the cost of algorithmic fidelity. By nature, these systems introduce a degree of off-policy learning, as data generated with model parameters from step $i$ may be used to train the model at step $i+1$ or beyond. Furthermore, asynchronous or partially asynchronous [he2025history, zhou2025april, gao2025rollpacker] RL systems often suffer from distributional skew, where faster-to-generate short samples disproportionately populate early training batches [meituan2025longcat]. These off-policy effects can diminish final model performance and complicate debugging and reproducibility [he2025nondeterminism]. Consequently, synchronous (or "on-policy") rollout remains critical in many settings for ensuring methodical evaluation, reproducibility, and strict adherence to the underlying algorithm’s assumptions. This paper, therefore, focuses on optimizing the synchronous case, although we note that our techniques can also be adapted to improve asynchronous settings.
@@ 2.2 Key Challenges in Rollout
To mitigate the long-tail effect and improve rollout scalability, recent works [zhong2025streamrl, fu2025areal, han2025asyncflow, sheng2025laminar, he2025history, zhou2025april, gao2025rollpacker] have proposed asynchronous RL. Unlike synchronous methods, where training experiences must all originate from the current policy, asynchronous methods allow training on stale rollout data from previous policy iterations, introducing off-policyness. By staging long-tail rollouts in advance or repacking requests, asynchronous RL can significantly reduce long-tail impact. However, to achieve the best convergence and maintain reward stability, many RL experiments still prefer synchronous training.
Speculative decoding (SD) offers a promising approach to accelerate the long-tail phase while preserving synchronous RL training guarantees. Existing SD techniques generally fall into two categories: model-based and model-free methods. Model-based SD approaches face key limitations in RL scenarios due to their static draft models and substantial overhead. First, the target LLM in RL is continuously updated, causing rapid "model drift". This divergence quickly renders the static draft model’s predictions obsolete, leading to a declining acceptance rate that nullifies potential performance gains. Second, the draft model itself introduces non-negligible latency and GPU memory consumption. In rollout scenarios characterized by large-batch offline inference workloads, such overhead becomes prohibitive. For model-free SD methods, n-gram-based approaches [fu2024break, vllm2025] are commonly employed. However, n-gram techniques require matching against the entire input sequence when generating draft tokens, incurring substantial overhead in long-context scenarios. Furthermore, n-gram methods struggle to aggregate multiple potential matches with low computational complexity, resulting in suboptimal draft token generation. To address these limitations of existing SD approaches, we propose grouped speculative decoding based on compressed suffix trees [weiner1973linear]. As a model-free method, it significantly reduces draft complexity while leveraging pattern similarity among requests within the same group to generate high-quality draft tokens. The detailed design of our SD approach is presented in §4.4.
@@ 7 Related Work
On-Policy RL Optimization. Some efforts have been made to optimize the RL workflow at both the system and algorithmic levels. RealHF [mei2024real] dynamically reallocates model parameters and memory budgets during RLHF to optimize utilization. RLHFuse [zhong2025optimizing] proposes stage fusion to overlap reward and experience computation with the rollout long tail, thereby improving resource utilization. RLBoost [wu2025rlboost] utilizes preemptible fragmented GPU resources to accelerate rollout at low cost. Nevertheless, these approaches do not adequately address the long-tail latency problem. For the long-tail latency problem in decoding, speculative decoding [leviathan2023fast] (SD) is regarded as an effective optimization technique. A series of works [li2024eagle1, li2024eagle2, li2025eagle3, cai2024medusa, liu2024deepseekv3] employ draft models or modules to generate draft tokens for target models, thereby reducing per-request latency. Despite achieving substantial speedup in small-batch scenarios, the effectiveness of these model-based speculative decoding algorithms is limited in RL scenarios, where model drift and the substantial overhead under large batch sizes result in significant performance degradation. Other works [fu2024break, oliaro2025suffixdecoding, he2025history, liu2025spec] have proposed model-free SD methods. RhymeRL [he2025history] and SPEC-RL [liu2025spec] propose utilizing historical sequences generated from previous RL iterations as references to accelerate decoding. However, in most state-of-the-art model rollout scenarios, different iterations sample different data to improve generalization, leaving no historical sequences to reuse. In contrast, Seer exploits the similarity among requests within the same group, employing online context learning within a single iteration to enhance speculative decoding effectiveness.
Off-Policy RL Optimization. Recently, many works [zhong2025streamrl, fu2025areal, han2025asyncflow, sheng2025laminar, he2025history, zhou2025april, gao2025rollpacker] have proposed RL workflows that do not maintain complete logical synchronization, trading a degree of off-policy behavior for improved RL efficiency. RhymeRL [he2025history], APRIL [zhou2025april], and RollPacker [gao2025rollpacker] reduce rollout tail latency through tail clipping and request repacking techniques. StreamRL [zhong2025streamrl], Areal [fu2025areal], and Laminar [sheng2025laminar] achieve complete decoupling of training and inference, substantially improving training concurrency and resource utilization. While asynchronous training improves RL efficiency, it introduces risks of instability and accuracy degradation, limiting its applicability in research and production environments where algorithmic fidelity is critical. Seer significantly reduces rollout tail latency through fine-grained scheduling of online context learning within the rollout phase, while maintaining consistency with on-policy algorithms.

### v2

@@ 1 Introduction
To improve hardware utilization, recent works have explored asynchronous rollout systems [zhong2025streamrl, fu2025areal, han2025asyncflow, sheng2025laminar, he2025history, piche2025pipelinerl], which overlap the rollout and training phases. While this approach can reduce end-to-end iteration time, it comes at the cost of algorithmic fidelity. By nature, these systems introduce a degree of off-policy learning, as data generated with model parameters from step $i$ may be used to train the model at step $i+1$ or beyond. Furthermore, asynchronous or non-strictly synchronous [zhou2025april, gao2025rollpacker] RL systems often suffer from distributional skew, where faster-to-generate short samples disproportionately populate early training batches [meituan2025longcat]. These off-policy effects can diminish final model performance and complicate debugging and reproducibility [he2025nondeterminism]. Consequently, synchronous (or "on-policy") rollout remains critical in many settings for ensuring methodical evaluation, reproducibility, and strict adherence to the underlying algorithm’s assumptions. This paper, therefore, focuses on optimizing the synchronous case, although we note that our techniques can also be adapted to improve asynchronous settings.
@@ 2.2 Key Challenges in Rollout
To mitigate the long-tail effect and improve rollout scalability, recent works [zhong2025streamrl, fu2025areal, han2025asyncflow, sheng2025laminar, he2025history] have proposed asynchronous RL. Unlike synchronous methods, where training experiences must all originate from the current policy, asynchronous methods allow training on stale rollout data from previous policy iterations, introducing off-policyness. Other works [zhou2025april, gao2025rollpacker], while maintaining synchrony, allow deferring a portion of requests to subsequent iterations, which cannot achieve iteration-level consistency compared to strict synchronous RL systems. By tail clipping or repacking requests, asynchronous or non-strictly synchronous RL can significantly reduce long-tail impact. However, to achieve the best convergence and maintain reward stability, many RL experiments still prefer synchronous training.
Speculative decoding (SD) offers a promising approach to accelerate the long-tail phase while preserving synchronous RL training guarantees. Existing SD techniques generally fall into two categories: model-based and model-free methods. Model-based SD approaches face key limitations in RL scenarios due to their static draft models and substantial overhead. First, the target LLM in RL is continuously updated, causing rapid "model drift". This divergence quickly renders the static draft model’s predictions obsolete, leading to a declining acceptance rate that nullifies potential performance gains. Second, the draft model itself introduces non-negligible latency and GPU memory consumption. In rollout scenarios characterized by large-batch offline inference workloads, such overhead becomes prohibitive. For model-free SD methods, n-gram-based approaches [fu2024break, vllm2025] are commonly employed. However, n-gram techniques require matching against the entire input sequence when generating draft tokens, incurring substantial overhead in long-context scenarios. Furthermore, n-gram methods struggle to aggregate multiple potential matches with low computational complexity, resulting in suboptimal draft token generation. To address these limitations of existing SD approaches, we propose grouped speculative decoding based on compressed suffix trees [weiner1973linear, hu2025sam, oliaro2025suffixdecoding]. As a model-free method, it significantly reduces draft complexity while leveraging pattern similarity among requests within the same group to generate high-quality draft tokens. The detailed design of our SD approach is presented in §4.4.
@@ 7 Related Work
On-Policy RL Optimization. Some efforts have been made to optimize the RL workflow at both the system and algorithmic levels. RealHF [mei2024real] dynamically reallocates model parameters and memory budgets during RLHF to optimize utilization. RLHFuse [zhong2025optimizing] proposes stage fusion to overlap reward and experience computation with the rollout long tail, thereby improving resource utilization. RLBoost [wu2025rlboost] utilizes preemptible fragmented GPU resources to accelerate rollout at low cost. Nevertheless, these approaches do not adequately address the long-tail latency problem. For the long-tail latency problem in decoding, speculative decoding [leviathan2023fast] (SD) is regarded as an effective optimization technique. A series of works [li2024eagle1, li2024eagle2, li2025eagle3, cai2024medusa, liu2024deepseekv3] employ draft models or modules to generate draft tokens for target models, thereby reducing per-request latency. Despite achieving substantial speedup in small-batch scenarios, the effectiveness of these model-based speculative decoding algorithms is limited in RL scenarios, where model drift and the substantial overhead under large batch sizes result in significant performance degradation. Other works [fu2024break, oliaro2025suffixdecoding, hu2025sam, he2025history, liu2025spec] have proposed model-free SD methods. RhymeRL [he2025history] and SPEC-RL [liu2025spec] propose utilizing historical sequences generated from previous RL iterations as references to accelerate decoding. However, in most state-of-the-art model rollout scenarios, different iterations sample different data to improve generalization, leaving no historical sequences to reuse. In contrast, Seer exploits the similarity among requests within the same group, employing online context learning within a single iteration to enhance speculative decoding effectiveness.
Off-Policy RL Optimization. Recently, many works [zhong2025streamrl, fu2025areal, han2025asyncflow, sheng2025laminar, he2025history, zhou2025april, gao2025rollpacker, piche2025pipelinerl, lu2025part] have proposed RL workflows that do not maintain complete logical synchronization, trading a degree of off-policy behavior for improved RL efficiency. RhymeRL [he2025history], APRIL [zhou2025april], and RollPacker [gao2025rollpacker] reduce rollout tail latency through tail clipping and request repacking techniques. StreamRL [zhong2025streamrl], Areal [fu2025areal], and Laminar [sheng2025laminar] achieve complete decoupling of training and inference, substantially improving training concurrency and resource utilization. While asynchronous training improves RL efficiency, it introduces risks of instability and accuracy degradation, limiting its applicability in research and production environments where algorithmic fidelity is critical. Seer significantly reduces rollout tail latency through fine-grained scheduling of online context learning within the rollout phase, while maintaining consistency with on-policy algorithms.